I came across the following question today:
A rectangular parcel of length x metres, width k metres and height k metres is to be sent through the post. The total length and girth of the parcel is to be exactly 2 metres. Show that the volume of the parcel is
V = x/16(2-x)^2
What I do not understand is, is the x in the volume equation the length of the parcel (ie 2m) or a different variable?
Also, does (2-x)^2 mean
(2-x)(2-x) ie 4-4x+x^2?

Found 2 solutions by ikleyn, math_tutor2020:
Answer by ikleyn(52777)

.
I came across the following question today:
A rectangular parcel of length x metres, width k metres and height k metres is to be sent through the post. The total length and girth of the parcel is to be exactly 2 metres. Show that the volume of the parcel is
V = x/16(2-x)^2
What I do not understand is, is the x in the volume equation the length of the parcel (ie 2m) or a different variable?
Also, does (2-x)^2 mean
(2-x)(2-x) ie 4-4x+x^2?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~


As I read this post,  I see that this assignment is  ILLOGICAL  and is as far
from to be a true  Math problem as the ground is far from heavens.

It has nothing in common with  Math problems,  as people traditionally understand this conception of  Math problems.


Regarding your last question,  whether does  (2-x)^2  mean  (2-x)(2-x)  ie  4-4x+x^2,
the answer is  " Yes ".



Answer by math_tutor2020(3816)


x = length
k = width
k = height

volume of rectangular block = length*width*height
V = x*k*k
V = x*k^2

If length + width = 2 meters, then,
x+k = 2
k = 2-x

So,
V = x*k^2
V = x*(2-x)^2
V = x*(2-x)(2-x)
V = x*(4-4x+x^2)
V = x^3-4x^2+4x

I don't know where your textbook got the x/16 from.